# In Defence of Post-hoc Explainability

### v1

@@ 4.1.1 Mediated Understanding
Scientific understanding through machine learning emerges not through direct model interpretation, but through a complex process of mediated interaction. The concept of "mediated understanding" in CI describes how scientific knowledge emerges through the structured interaction between four key elements: model behaviour, interpretability methods, domain knowledge, and empirical validation. This principle recognises that scientific understanding through ML is inherently mediated – direct access to model mechanics is not necessary for scientific insight (Sullivan, 2022; Beisbart and Räz, 2022). Instead, understanding emerges through a dialectical process of interpretation, where the relationship between model understanding and phenomenal understanding is reciprocally constitutive.
The epistemic validity of post-hoc methods in CI stems from their role as structured mediators in a bidirectional knowledge-creation process. In one direction (Model → Phenomenon), interpretability methods reveal patterns in model behaviour, which then tentatively suggest hypotheses about phenomena. These hypotheses, when tested empirically, provide new phenomenal understanding. In the other direction (Phenomenon → Model), domain knowledge guides the selection and refinement of interpretability methods, while empirical validation helps refine our interpretive approaches and identify relevant model behaviours for investigation. This bidirectional mediation provides epistemic justification because it ensures that interpretability methods are not merely describing model behaviour, but are actively participating in a cycle of hypothesis generation, empirical validation, and knowledge refinement – the very essence of scientific inquiry.
Consider a medical diagnosis ML system as an illustrative example (Sullivan, 2022). Feature attribution methods serve as epistemic mediators by translating opaque model computations into testable hypotheses about biological mechanisms. When these model-derived insights are empirically validated against existing scientific evidence, they contribute to medical knowledge not despite their post-hoc nature, but precisely because their mediated interpretation enables systematic comparison between model behaviour and real-world phenomena. This demonstrates how CI’s concept of mediated understanding resolves the apparent conflict between post-hoc interpretation and scientific validity – the very process of mediation, when properly structured and validated, becomes a legitimate source of scientific knowledge. A detailed examination of how mediated understanding operates differently across theory-rich versus theory-poor contexts is provided in Appendix B.
@@ 4.1.2 Bounded Factivity
Building on our discussion of mediated understanding, which shows how scientific knowledge emerges through structured interactions between models and methods, we now turn to another key concept that supports CI’s defense of post-hoc interpretability: bounded factivity. This concept helps resolve fundamental tensions between the approximative nature of post-hoc methods and their epistemic value for scientific understanding.
In philosophy of science, the relationship between factivity and scientific understanding has been extensively debated. Traditional accounts often assume that genuine understanding requires strictly factual, true beliefs. However, recent work in philosophy of science has highlighted how idealisation and approximation play essential roles in scientific practice (Beisbart and Räz, 2022; Sullivan, 2022; Freiesleben et al., 2024). Scientists routinely use simplified models that deliberately deviate from reality to gain understanding of complex phenomena. These non-factive elements, rather than being mere compromises, often prove essential for scientific progress.
@@
This recognition of strategic simplification’s role in science helps us reconceptualise the epistemic status of post-hoc interpretability methods. Rather than demanding complete factivity – perfect correspondence between interpretation and model mechanics – CI advocates for what we term "bounded factivity": truth within explicitly acknowledged limits and simplifications. Recent empirical work underscores the importance of this bounded approach. While Bilodeau et al. (2024) demonstrated that many popular post-hoc methods perform no better than random attribution, the authors also showed that carefully designed, task-specific approaches can provide reliable insights. By aligning interpretability methods with specific scientific goals (Freiesleben et al., 2024) and validating them through systematic empirical testing, we can achieve meaningful understanding within acknowledged bounds, just as traditional scientific models advance understanding despite their simplifications.
The concept of bounded factivity finds a natural parallel in Herbert Simon’s bounded rationality (Simon, 1957; Wheeler, 2018) – both frameworks acknowledge inherent limitations while affirming the validity of strategic simplification. Just as bounded rationality accepts "satisficing" solutions under cognitive constraints, bounded factivity embraces carefully bounded approximations in interpretation. This pragmatic orientation helps justify post-hoc methods by showing how they can advance scientific understanding even without achieving perfect fidelity.
These parallels have important philosophical implications for how we justify post-hoc interpretability in scientific ML. Just as bounded rationality challenged idealised notions of human decision-making while preserving the validity of human judgment, bounded factivity challenges the assumption that perfect explanations are necessary for genuine scientific understanding while defending the epistemic value of post-hoc interpretations. It suggests that valid scientific understanding emerges through the careful management of trade-offs between accuracy and comprehensibility, realised through the process of mediated understanding where model behaviour, interpretive methods, and empirical validation interact cyclically. This mediated process, operating within explicitly acknowledged bounds, enables post-hoc methods to generate reliable scientific insights by systematically refining both our interpretation of models and our understanding of phenomena – precisely what well-designed post-hoc methods aim to achieve.
4.2 Re-assessing Criticisms of Post-hoc Models
4.2.1 Approximation and Fidelity
@@
Rudin (2019) and Ghassemi et al. (2021) argue that post-hoc explanations are problematic due to their approximative nature, this critique necessitates careful examination of distinct but related concerns: the factivity of explanations and the nature of understanding in scientific practice. The dilemma presents itself thus: completely accurate explanations of complex ML models would merely duplicate their opacity, while simplified explanations necessarily introduce some degree of falsehood. This apparent tension can be productively addressed through the lens of non-factive understanding in science (Beisbart and Räz, 2022).
@@
Interpretability exists on a spectrum, with increased epistemic value and practical utility correlating with higher degrees of interpretability. This approach aligns with the concept of verisimilitude, where approximations to truth, though imperfect, retain epistemic worth (Oddie, 2001). Although post-hoc explanations lack performance guarantees and do not fully capture model behaviour, this limitation need not compromise their epistemological value if we maintain awareness of departure (Kvanvig, 2009) – conscious recognition of where and how our explanations diverge from ground truth. Many scientific and analytical tools rely on strategic idealisations that, despite their non-factive nature, provide valuable insights and practical utility. The key is maintaining empirical accountability through testable predictions, situating approximations within relevant theoretical frameworks, and providing clear scope conditions for the validity of interpretations.
This perspective aligns with CI’s broader commitment to epistemic accessibility while acknowledging that accessibility often requires trade-offs with complete accuracy. Just as scientific models generally involve idealisations that technically violate factivity without compromising their utility for understanding, post-hoc explanations can provide genuine scientific insight even while containing strategic simplifications. This suggests that the key to maintaining scientific rigour lies not in perfect factivity, but in transparent acknowledgment of simplifications coupled with continuous refinement through empirical validation.
@@
The relationship between model, interpretation, and reality can be illuminated through Korzybski’s dictum, "The map is not the territory". Just as a map is a simplified representation of reality, both intrinsically interpretable models and post-hoc explanations are simplifications of the complex systems they represent. Accepting an intrinsically interpretable model as "understandable" and having some fidelity to the real world is philosophically analogous to accepting a post-hoc explanation that is "understandable" and has some fidelity to the original model. The fidelity between complex systems (real world or AI) and any model (intrinsic or post-hoc) is inherently imperfect, yet this imperfection does not negate their scientific value when properly bounded and validated.
4.2.2 Faithful Explanation and Confirmation Bias
Rudin (2019) identifies two potential pitfalls of post-hoc models – incomplete (local) explanations and unjustifiable explanations. The critique of local explanations underestimates their unique epistemological value in scientific practice. Rather than viewing local explanations as merely incomplete versions of global understanding, CI recognises them as distinct epistemic tools that offer granular insights into model behaviour. Through mediated understanding, these local insights can generate testable hypotheses about both model behaviour and phenomenal relationships, identify edge cases that reveal important patterns, and expose nuances that global explanations might miss. When properly bounded and validated, local explanations complement rather than compete with global understanding.
@@ 4.2.2 Faithful Explanation and Confirmation Bias
Regarding unjustifiable explanations, CI posits that even apparently problematic model behaviours – such as scientifically unsound judgments or confounding variables – can advance scientific understanding when properly interpreted. Through bounded factivity, we recognise that identifying flaws in model reasoning contributes valuable knowledge about both model limitations and phenomenal complexity. This aligns with how sciences historically progress through understanding both positive and negative results.
Ghassemi et al. (2021) raise a complementary concern about confirmation bias in interpreting post-hoc explanations, suggesting that humans might draw overconfident conclusions from potentially unreliable interpretations. This "interpretability gap" could potentially foster false confidence in the model’s reliability or fairness. The limitations Ghassemi et al. describe are not unique to AI explanations but are inherent in complex judgements, whether human or artificial.
@@
Human experts, like AI systems, can fall prey to confirmation bias, potentially leading to overconfidence in their interpretations or explanations. Following this reasoning to its logical conclusion, one might argue that we should be equally sceptical of human expert explanations as we are of AI-generated ones. Taken to an extreme, this line of thinking could lead to an argument for minimising reliance on expert explanations altogether, whether human or AI-generated. Instead, one might advocate for sole reliance on predefined, explicit "if-then" rules, aiming to eliminate the subjectivity and potential biases inherent in both human and AI interpretations. Yet, this conclusion overlooks the value of both human and AI-generated post-hoc explanations.
@@
Rather than suggesting we should abandon post-hoc explanations in favor of purely rule-based approaches, CI advocates for their refinement and systematic validation. Just as sciences have developed methods for managing human cognitive biases while preserving the value of expert insight, we can develop approaches to post-hoc interpretation that acknowledge limitations while maximising epistemic value. This calls the field’s attention to validation procedures and explicit acknowledgment of bounds.
5 Conclusion
Attempts to interpret AI models, particularly through post-hoc explanations, are analogous to human experts translating their intuitive impressions into deliberate explanations. Just as human experts often provide post-hoc rationalisations for their judgments, post-hoc explainability methods seek to make sense of the complex, compiled knowledge within AI. This process is inherently imperfect, as human experts also rarely access the full breadth of their internalised chunks and productions (Newell and Simon, 1972; Simon and Chase, 1973; Gobet and Clarkson, 2004). However, this parallel reveals instructive differences in how we can approach AI interpretation. Unlike the black box of human cognition, AI systems permit systematic investigation of their internal states, allowing us to precisely specify the bounds of factivity and empirically verify our approximations. This accessibility enables a more rigorous approach to post-hoc interpretation than is possible with human expertise.
The epistemological framework of Computational Interpretabilism suggests that post-hoc interpretability methods serve a crucial mediating function in scientific ML, analogous to how expert explanations bridge specialised knowledge and broader understanding. When medical experts translate complex diagnoses for patients, they necessarily simplify their understanding into comprehensible explanations. Similarly, post-hoc interpretability methods transform opaque model computations into scientifically meaningful insights. The epistemic value lies not in perfect mechanical reproduction of the underlying process, but in reliable knowledge generation that can be validated against empirical evidence and integrated with domain expertise.
Appendix A Functional Relationships in Interpretable and Post-hoc AI Models
Figure 1: A comparative analysis of intrinsically interpretable and post-hoc explainable AI models.
This diagram illustrates the key functional relationships in both intrinsically interpretable and post-hoc explainable AI models within scientific contexts. The framework is grounded in three fundamental functions: the true phenomenon $f(X)$, the learned model $h^{*}(X)$, and for post-hoc methods, the interpretable approximation $p^{*}(X)$.
In both cases, we begin with a true function $f(X)$ representing the natural phenomenon of interest. The model learns an approximation $h^{*}(X)$ within a hypothesis space $\mathcal{H}$, where $\mathcal{H}$ represents the set of possible model behaviours $\{h_{1},...,h_{n}\}$. For this analysis, we assume ideal conditions where the training data accurately represents the population distribution, allowing us to focus on the core relationships between these functions.
The key distinction between intrinsically interpretable and post-hoc approaches lies in how they achieve human comprehensibility:
  • In intrinsically interpretable models, $h^{*}(X)$ is constrained to be directly comprehensible, as indicated by the dashed boundary in the left diagram.
  • In post-hoc models, an additional function $p^{*}(X)$ is learned from a space of possible explanations $\mathcal{P}$ to approximate $h^{*}(X)$’s behaviour in an interpretable way.
The learning paths (shown in red for intrinsic and blue for post-hoc) illustrate how each approach navigates the tension between model capability and human comprehensibility. While intrinsically interpretable models maintain a direct connection to human understanding, post-hoc methods introduce a secondary layer of interpretation that can potentially capture more complex relationships between the model and the underlying phenomenon.
Appendix B Theoretical Contexts and Epistemic Value in Machine Learning Interpretation
While Andrews [2023] demonstrates how theoretical considerations play essential roles throughout the ML pipeline in cases like AlphaFold – from data engineering to architecture design and model evaluation – not all AI models for science are built with this same level of theoretical rigour. CI expands this understanding by demonstrating that post-hoc interpretability methods can be epistemically justified across the full spectrum of theoretical contexts through structured mediated understanding. This broadens the scope of scientific ML beyond just theory-rich applications.
In theory-rich cases like AlphaFold, CI shows how post-hoc interpretability methods [Tan and Zhang, 2023] derive their epistemic validity from their role in a theoretically-grounded mediation process. They help validate whether models implement intended theoretical principles, enable scientists to verify if learned representations align with established theories, support discovery of potential new theoretical insights, and provide means to examine whether theoretical assumptions embedded in model design function as intended. The mediation process here acts as a bridge between explicit theoretical commitments and model behaviour, with interpretability methods serving as epistemically-justified tools for both validation and discovery precisely because they operate within a well-defined theoretical framework.
Even more significantly, CI demonstrates how post-hoc methods maintain epistemic validity in theory-poor contexts where ML models are applied with minimal theoretical consideration. Here, the mediation process takes on a different but equally valid epistemic role: helping uncover implicit theoretical assumptions embedded in data selection and preprocessing, providing ways to retrospectively examine model behaviour against domain knowledge, revealing potential biases or problematic patterns, and identifying opportunities for theoretical incorporation. Through mediated understanding, these methods create an epistemically rigorous path to reconstruct and validate theoretical frameworks that might have been overlooked in model development – effectively building theoretical bridges where none previously existed.
CI thus resolves a key epistemological challenge in machine learning interpretation: how to justify post-hoc interpretability methods across varying levels of theoretical grounding. The framework demonstrates that through properly structured mediation processes, these methods maintain epistemic validity whether they are validating existing theoretical frameworks or helping construct new ones. In both cases, post-hoc interpretability methods hold epistemic value because they enable systematic scientific scrutiny of how these models operate within their domains, regardless of how deliberately theory was incorporated into their development. This unified treatment of theory-rich and theory-poor contexts represents a significant advance in our understanding of how ML can contribute to scientific knowledge across diverse applications.

### v2

@@ 4.1.1 Mediated Understanding
Scientific understanding through machine learning emerges not through direct model interpretation, but through a complex process of mediated interaction. The concept of mediated understanding describes how scientific knowledge emerges through the structured interaction between four key elements: model behaviour, interpretability methods, domain knowledge, and empirical validation. This principle recognises that scientific understanding through ML is inherently mediated – where direct access to model mechanics is neither necessary nor sufficient for scientific insight Sullivan (2022); Beisbart and Räz (2022).
The epistemic validity of post-hoc methods stems from their role as mediators in a bidirectional knowledge-creation process. In one direction (Model → Phenomenon), interpretability methods reveal patterns in model behaviour, which then tentatively suggest hypotheses about phenomena. These hypotheses, when tested empirically, provide new phenomenal understanding. In the other direction (Phenomenon → Model), domain knowledge guides the selection and refinement of interpretability methods, whilst empirical validation helps refine our interpretive approaches and identify relevant model behaviours for investigation. This bidirectional mediation provides epistemic justification because it ensures that interpretability methods are not merely describing model behaviour, but are actively participating in a cycle of hypothesis generation, empirical validation, and knowledge refinement – the very essence of scientific inquiry.
@@ 4.1.2 Bounded Factivity
Building on our discussion of mediated understanding, we now turn to bounded factivity, which helps resolve fundamental tensions between the approximative nature of post-hoc methods and their epistemic value for scientific understanding. Rather than demanding complete factivity – perfect correspondence between interpretation and model mechanics – bounded factivity acknowledges truth within explicitly acknowledged limits and simplifications. This aligns with traditional scientific practice, where scientists routinely use simplified models that deliberately deviate from reality to gain understanding of complex phenomena Beisbart and Räz (2022); Sullivan (2022).
@@
The recognition of strategic simplification’s role in science helps reconceptualise the epistemic status of post-hoc interpretability methods. Recent empirical work by Bilodeau et al. (2024)] demonstrates that whilst many general-purpose post-hoc methods may be unreliable, carefully designed, task-specific approaches can provide reliable insights within bounded contexts. By aligning interpretability methods with specific scientific goals Freiesleben et al. (2024) and validating them through systematic empirical testing Sullivan (2022), we can achieve meaningful understanding within acknowledged bounds, just as traditional scientific models advance understanding despite their simplifications.
4.1.3 Structured Process of Scientific Discovery
The philosophical framework we have established can be operationalised into a structured process for generating scientific knowledge through post-hoc methods (2, LABEL:app). This process, illustrated through a cyclic interaction between model behaviour, interpretability methods, and domain knowledge, demonstrates how post-hoc methods can contribute to scientific understanding whilst maintaining epistemic rigour. The diagram pairs each theoretical component with a concrete example from type 2 diabetes research (detailed in the next subsection), providing a practical roadmap for how post-hoc interpretability methods can generate legitimate scientific knowledge.
4.2 Bridging Philosophy and Practice: A Case Study
The recent work by Klein et al. (2024) on explainable AI-based analysis of pancreatic tissue in type 2 diabetes (T2D) provides a compelling demonstration of our philosophical framework in action. Their study exemplifies how post-hoc explainability methods can generate legitimate scientific knowledge through systematic application of mediated understanding and bounded factivity principles.
The study demonstrates mediated understanding through carefully orchestrated interactions between model behaviour, interpretability methods, and domain expertise. The translation phase employed multiple XAI techniques – attention mechanisms and attribution methods —- to convert complex model computations into interpretable visualisations. This translation was inherently theory-laden: the selection of six immunostaining markers (insulin, glucagon, somatostatin, PECAM1, perilipin 1, and tubulin beta 3) reflected theoretical understanding whilst remaining open to novel discoveries. Remarkably, their best-performing model focused on islet $\alpha$- and $\delta$-cells along with neuronal structures rather than $\beta$-cells, challenging traditional assumptions about T2D pathophysiology. This unexpected finding emerged through the interaction between model behaviour and domain expertise, demonstrating how post-hoc methods can reveal insights that transcend existing theoretical frameworks.
The bounded factivity principle manifests powerfully in their validation approach. Rather than claiming complete mechanistic understanding, the researchers demonstrated scientific rigour – the strict application of the scientific method to experimental design and interpretation (Klein et al., 2024; Weber et al., 2024) – by explicitly bounding their claims within testable hypotheses. They developed targeted biomarkers to quantify specific observations: adipocyte-to-islet proximity (measured in micrometres), fibrosis density patterns, and morphological features. These biomarkers underwent systematic validation through generalised mixed linear models (GMLM), controlling for clinical covariates (age, sex, BMI) and random effects (cohort, staining method). Notably, they achieved statistical significance for key associations, whilst acknowledging limitations such as segmentation challenges.
Klein et al. (2024) compellingly demonstrated how post-hoc explainability methods advance scientific understanding beyond mere model interpretation. Their discovery that islets in T2D patients are significantly closer to adipocytes suggests paracrine effects impacting islet function – a hypothesis that emerged directly from XAI-guided analysis. The unexpected finding that $\alpha$-cells, $\delta$-cells, and neuronal markers outperformed $\beta$-cell markers in prediction challenges the $\beta$-cell-centric view of T2D and suggests more complex pathophysiology involving islet innervation. These insights were not merely descriptions of model behaviour but genuine contributions to understanding T2D pathogenesis, validated through statistical analysis and grounded in biological plausibility. The study thus exemplifies how post-hoc methods, when embedded within structured scientific practices and constrained by bounded factivity, serve as legitimate tools for scientific discovery.
5 Re-assessing Alternative Views
5.1 Approximation and Fidelity
@@
Rudin (2019) and Ghassemi et al. (2021) argue that post-hoc explanations are problematic due to their approximative nature. This critique necessitates careful examination of distinct but related concerns: the factivity of explanations and the nature of understanding in scientific practice. The dilemma presents itself thus, completely accurate explanations of complex ML models would merely duplicate their opacity, whilst simplified explanations necessarily introduce some degree of falsehood. This apparent tension can be productively addressed through the lens of non-factive understanding in science Beisbart and Räz (2022).
@@
Interpretability exists on a spectrum, with increased epistemic value and practical utility correlating with higher degrees of interpretability. This approach aligns with the concept of verisimilitude (“closeness to truth”), where approximations to truth, though imperfect, retain epistemic worth Oddie (2001). Although post-hoc explanations lack performance guarantees and do not fully capture model behaviour, this limitation need not compromise their epistemological value if we maintain awareness of departure Kvanvig (2009) – conscious recognition of where and how our explanations diverge from ground truth. Many scientific and analytical tools rely on strategic idealisations that, despite their non-factive nature, provide valuable insights and practical utility. The key is maintaining empirical accountability through testable predictions, situating approximations within relevant theoretical frameworks, and providing clear scope conditions for the validity of interpretations.
@@
Just as scientific models generally involve idealisations that technically violate factivity without compromising their utility for understanding, post-hoc explanations can provide genuine scientific insight even while containing strategic simplifications. This suggests that the key to maintaining scientific rigour lies not in perfect factivity, but in transparent acknowledgment of simplifications coupled with continuous refinement through empirical validation. Both intrinsically interpretable models and post-hoc explanations are simplifications of the complex systems they represent. Accepting an intrinsically interpretable model as ‘understandable’ and having some fidelity to the real world is philosophically analogous to accepting a post-hoc explanation that is ‘understandable’ and has some fidelity to the original model. The fidelity between complex systems (real world or AI) and any model (intrinsic or post-hoc) is inherently imperfect, yet this imperfection does not negate their scientific value when properly bounded and validated.
5.2 Faithful Explanation and Confirmation Bias
@@
Rudin’s [2019] critique of post-hoc methods focuses on two potential pitfalls: incomplete (local) explanations and unjustifiable explanations. However, local explanations could serve as distinct epistemic tools that offer granular insights into model behaviour. These local insights can generate testable hypotheses about both model behaviour and phenomenal relationships, identify edge cases that reveal important patterns, and expose nuances that global explanations might miss. Local explanations may complement rather than compete with global understanding. Regarding unjustifiable explanations, our framework suggests that even apparently problematic model behaviours – such as scientifically unsound judgments or confounding variables – can advance scientific understanding when properly interpreted. We recognise that identifying flaws in model reasoning contributes valuable knowledge about both model limitations and phenomenal complexity. This aligns with how sciences historically progress through understanding both positive and negative results.
@@
Ghassemi et al. (2021) raise a complementary concern about confirmation bias in interpreting post-hoc explanations, suggesting that humans might draw overconfident conclusions from potentially unreliable interpretations. This "interpretability gap" could potentially foster false confidence in the model’s reliability or fairness. The limitations Ghassemi et al. describe are not unique to AI explanations but are inherent in complex judgements, whether human or artificial. Human experts, like AI systems, can fall prey to confirmation bias, potentially leading to overconfidence in their interpretations or explanations. Following this reasoning to its logical conclusion, one might argue that we should be equally sceptical of human expert explanations as we are of AI-generated ones. Taken to an extreme, this line of thinking could lead to an argument for minimising reliance on expert explanations altogether, whether human or AI-generated. Instead, one might advocate for sole reliance on predefined, explicit "if-then" rules, aiming to eliminate the subjectivity and potential biases inherent in both human and AI interpretations. Yet, this conclusion overlooks the value of both human and AI-generated post-hoc explanations.
@@
Rather than suggesting we should abandon post-hoc explanations in favour of purely rule-based approaches, our framework advocates for their refinement and systematic validation. Just as sciences have developed methods for managing human cognitive biases while preserving the value of expert insight, we can develop approaches to post-hoc interpretation that acknowledge limitations while maximising epistemic value through careful bounds and empirical validation.
6 Discussion
This position paper has several limitations that warrant acknowledgement. First, our philosophical framework relies primarily on a single empirical case study (Klein et al., 2024) to demonstrate practical application. Whilst this study provides some evidence, additional examples across diverse scientific domains would strengthen our argument. Second, we focus exclusively on supervised learning for scientific discovery, excluding important ML paradigms such as unsupervised learning, reinforcement learning, and generative models, which may require different epistemological considerations. Third, we deliberately focus on epistemic rather than normative considerations, leaving important questions about the ethical obligations and social responsibilities of XAI unexplored (see Appendices C.6 and C.5 for relevant normative discussions). Finally, whilst we introduce the concept of "bounded factivity," we remain at the philosophical level without operationalising it into measurable criteria. Though Wei et al. (2024) demonstrate fine-grained metrics for evaluating post-hoc method robustness, translating such approaches to scientific contexts remains unclear (add Friesleban). What constitutes sufficient "bounds" for scientific practice? How do we measure faithfulness when the goal is phenomenal understanding rather than model fidelity? Our focus was establishing how post-hoc methods can be philosophically justified within scientific practices; developing empirical measures and concrete desiderata informed by philosophy of science represents an important but separate endeavour for future work.
Attempts to interpret AI models, particularly through post-hoc methods, bear a meaningful resemblance to the way human experts articulate their intuitive judgements. Much like post-hoc rationalisations offered by human specialists – who often cannot access the full breadth of their internalised cognitive processes—these AI explanations seek to render comprehensible the dense, layered computations embedded in modern models (Newell and Simon, 1972; Simon and Chase, 1973; Gobet and Clarkson, 2004). This parallel highlights not only the approximative nature of both forms of explanation but also a critical distinction: unlike the human mind, AI systems permit systematic interrogation of their internal states. This accessibility allows for more transparent bounds on factivity and the possibility of empirical validation—advantages that position post-hoc explainability as a potentially more rigorous and inspectable form of interpretation than human expertise affords. In this view, post-hoc methods are not peripheral conveniences but central instruments in the pursuit of scientific insight from ML systems. While their approximative nature imposes limits, these limits are not disqualifying – provided they are acknowledged, bounded, and subjected to empirical scrutiny. Post-hoc explainability, when properly constrained and philosophically grounded, can be an effective conduit for phenomenal understanding within scientific practice.
Appendix A Conceptual Comparison of Intrinsic and Post-hoc Interpretability Models
(a) Intrinsically interpretable models constrain the hypothesis space $\mathcal{H}$ to human-comprehensible function classes (e.g., linear models, decision trees). The learned model $h^{*}(X)$ approximates the phenomenon $f(X)$ whilst remaining directly interpretable by design.
(b) Post-hoc explainability allows models to be learned from an unconstrained hypothesis space $\mathcal{H}$, potentially capturing complex patterns through $h^{*}(X)$. A separate explanation function $p^{*}(X)$ is then derived (e.g., via LIME, SHAP) to approximate and interpret $h^{*}(X)$’s behaviour in human-understandable terms.
Figure 1: Comparison of intrinsic and post-hoc interpretability models.
Appendix B Framework for Scientific Knowledge Generation through Post-hoc Methods
Figure 2: Framework for Scientific Knowledge Generation through Post-hoc Methods. The diagram illustrates the cyclic process of generating scientific understanding from ML models through post-hoc interpretability methods. Each component shows both the theoretical principle (dark grey boxes) and its practical application in T2D research (light grey boxes). The process begins with translation of model behaviour, progresses through method selection and hypothesis generation, and culminates in empirical validation through instrumental and theoretical refinement. Green elements represent the key stages in the knowledge generation pipeline. Arrows indicate the flow and interactions between components, demonstrating how post-hoc methods mediate between model behaviour and scientific understanding.